OpenStudy (anonymous):

A rope is square root of two hundred fifty units long. The rope is cut into two pieces, so that the lengths of the pieces are in the ratio 3:2 . What is the length of the longer piece expressed in simplest radical form? (1 point) 3square root of tenunits 10square root fiveunits 10 square root two units 2square root of ten units

OpenStudy (anonymous):

\[3x+2x = 5x = \sqrt{250} = 5\sqrt{10}\] To find x, \[x={5\sqrt{10} \over 5} = \sqrt{10} \] Since the longer piece is 3x, you take \[\sqrt{10} \times 3 =3\sqrt{10}\] does that make sense?

OpenStudy (anonymous):

Yes it does :) Thank you!!!